Bach Corp. had additions to retained earnings for the year just ended of $435,000. The firm
paid out $160,000 in cash dividends, and it has ending total equity of $5.3 million. If the
company currently has 210,000 shares of common stock outstanding, what are the earnings per
share? Dividends per share? Book value per share? If the stock currently sells for $63 per share,
what is the price-earnings ratio?
Answer #1

Earnings per share =(Dividend+Retained Earnings)/Number of Shares =(160000+435000)/210000=2.83

Dividend per Share = Cash Dividends/Number of shares =160000/210000=0.76

Book Value per Share =Equity/Number of Shares =5300000/210000=25.24

Price to Earnings Ratio =63/2.8333=22.24